OpenStudy (anonymous):

can anyone please help me on this question What are the first five terms of the pattern with the formula 2n + 2? The first five terms are , , , , and .

OpenStudy (anonymous):

if n=1, 2*1+2 = 4, n=2: 2*2+2=6, n=3: 2*3+2=8, n=4: 2*4+2=10, n=5: 2*5+2=12 so 2, 6, 8, 10, 12

OpenStudy (anonymous):

2, 6, 8, 10, 12 is the answer